# One-loop corrections to the neutrino in the N-B-LSSM

### v1

@@ I Introduction
Recent experimental data, such as results released in 2023 by the LHCb collaboration5lhc , have further propelled studies on neutrino masses and lepton flavor violation processes. These data not only provide precise measurements of neutrino mixing angles but also impose new constraints on the upper and lower bounds of neutrino masses. Such constraints are vital for unraveling the nature of neutrinos and their role in the universe. The present $3\sigma$ limits for the neutrino expriment data are6pdg
@@
$6.99\times 10^{-5}\leq\bigtriangleup{m^{2}}_{\odot}\leq 8.07\times 10^{-5}$
@@
Based on the MSSM, the minimal supersymmetric extension of the SM with local B-L norm symmetry (N-B-LSSM) is obtained by extending the gauge symmetry group to $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}\times U(1)_{B-L}$7h . In the N-B-LSSM, we have discussed certain two-loop contributions to muon anomalous magnetic dipole moment, lepton flavour-violating decays $l_{j}\rightarrow l_{i\gamma},l_{j}\rightarrow 3l_{i}$ and $\mu\rightarrow e+q\overline{q}$, but the neutrino mass problem has not yet been studied7h . In this paper, we study the neutrino mass correction problem under the N-B-LSSM, neutrinos acquire tiny masses at tree level through the seesaw mechanism8seesaw ; 9sm7 ; 102zlxzqqb , facilitated by the introduction of right-handed neutrino fields. However, one-loop corrections to the neutrino mass matrix are crucial for a comprehensive understanding of neutrino masses and mixings10yyl22 ; 11sm . These corrections arise from various sources, including virtual slepton-chargino, sneutrino-neutralino, and Higgs-charged lepton loops. The inclusion of one-loop effects is essential to match the precise experimental data on neutrino oscillations, which provide stringent constraints on the neutrino mass squared differences and mixing angles.
In ref12tff , the authors discuss the supersymmetric neutrino mass under the one-loop approximation using the on-shell renormalization scheme without taking into account lepton number conservation and R-parity. There are also some work related to the mass and mixing of neutrinos13yyl ; 101zlxz ; 14yyl33 ; 15tff8 . This paper focuses on the one-loop corrections to the neutrino mass matrix in the N-B-LSSM. We employ the method of mass eigenstates to calculate these corrections and analyze their impact on the neutrino mass spectrum and mixing angles. Our study aims to explore the parameter space of the N-B-LSSM that is consistent with the current neutrino oscillation data. By doing so, we aim to shed light on the underlying mechanisms of neutrino mass generation and to provide insights into the new physics beyond the SM.
@@ II The N-B-LSSM
Under this model, $Y_{u,d,e,\nu,\chi}$ in the superpotential denotes the Yukawa coupling coefficient. While $\lambda$, $\lambda_{2}$ and $\kappa$ are dimensionless couplings constant. $\hat{\chi}_{1},~{}\hat{\chi}_{2},~{}\hat{S}$ are three Higgs singlets.
@@
The specific explicit expressions of two Higgs doublets and three Higgs singlets are as follows
@@ III ANALYTICAL FORMULA
In this section, we compute the one-loop correction to the neutrino mass under the N-B-LSSM. The general form of the self-energy of ${\nu^{0}_{i}}-{\nu^{0}_{j}}$ is as follows
$S_{ij}(k)={a_{ij}m_{j}\omega_{-}}+{b_{ij}m_{i}\omega_{+}}+{c_{ij}{k\!\!\!/}\omega_{-}}+{d_{ij}{k\!\!\!/}\omega_{+}}.$ (31)
We can write $a_{ij},b_{ij},c_{ij}$ and $d_{ij}$ as expressions of $k^{2}$12tff ; 21tff12
@@
$S_{ij}^{REN}(k)u_{i}(k)|_{k^{2}=m_{i}^{2}}=0,$
$\bar{u}_{j}(k)S_{ij}^{REN}(k)|_{k^{2}=m_{j}^{2}}=0,$ (34)
which yield solutions
@@
$S_{ij}^{REN}(k)=({m_{i}}^{2}b_{ij}^{1}+{m_{i}}^{2}c_{ij}^{1}+{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+a_{ij}^{1}k^{2})m_{j}\omega_{-}$
$\hskip 42.67912pt+({m_{j}}^{2}a_{ij}^{1}+{m_{j}}^{2}c_{ij}^{1}+{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+b_{ij}^{1}k^{2})m_{i}\omega_{+}$
$\hskip 42.67912pt+\Big{(}-{m_{j}}^{2}a_{ij}^{1}-{m_{i}}^{2}b_{ij}-({m_{i}}^{2}+{m_{j}}^{2})c_{ij}^{1}-{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+b_{ij}^{1}k^{2}\Big{)}$
$\hskip 42.67912pt+\Big{(}-{m_{i}}^{2}{m_{j}}^{2}a_{ij}^{1}-{m_{i}}^{2}{m_{j}}^{2}b_{ij}^{1}-{m_{i}}^{2}{m_{j}}^{2}c_{ij}^{1}-({m_{i}}^{2}+{m_{j}}^{2})f_{ij}^{1}+d_{ij}^{1}k^{2}\Big{)}$
$\hskip 42.67912pt=(k\!\!\!/-m_{j})\hat{S}_{ij}(k)(k\!\!\!/-m_{i})$ (36)
@@
$\hat{S}_{ij}(k)=a_{ij}^{1}m_{j}\omega_{+}+b_{ij}^{1}m_{i}\omega_{-}+c_{ij}^{1}(m_{i}\omega_{-}+m_{j}\omega_{+}+k\!\!\!/\omega_{+})$
$\hskip 42.67912pt+d_{ij}^{1}(m_{i}\omega_{+}+m_{j}\omega_{+}+k\!\!\!/\omega_{-})$ (37)
@@
$\delta{m_{ij}^{R}}=({m_{j}}^{2}a_{ij}^{1}+{m_{j}}^{2}c_{ij}^{1}+{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+b_{ij}^{1}k^{2})m_{i}.$ (38)
@@
$=(\delta_{ij}+\delta Z_{ij}^{L})(k\!\!\!/-m_{i}^{tr}-\delta m_{ij}^{L}+\delta Z_{ij}^{L}m_{i}^{tr})\omega_{-}$
$+(\delta_{ij}+\delta Z_{ij}^{R})(k\!\!\!/-m_{j}^{tr}-\delta m_{ij}^{R}+\delta Z_{ij}^{R}m_{j}^{tr})\omega_{+},$ (39)
@@
$\delta m^{loop}_{ij}=3m_{i}^{tr}(m_{i}^{tr})^{2}a_{ij}^{1}+\Big{(}m_{i}^{tr}m_{j}^{tr}+(m_{i}^{tr})^{2}+(m_{j}^{tr})^{2}\Big{)}b_{ij}^{1}$
$\hskip 34.14322pt+\Big{(}(m_{i}^{tr})^{2}m_{j}^{tr}+3m_{i}^{tr}(m_{j}^{tr})^{2}\Big{)}c_{ij}^{1}+\Big{(}3(m_{i}^{tr})^{2}m_{j}^{tr}+m_{i}^{tr}(m_{i}^{tr})^{2}\Big{)}d_{ij}^{1}$ (40)
@@
The mass of neutrinos are obtained through the rotation matrix $Z_{N,\nu}$ using the formula $Z_{N,\nu}^{T}M_{N}Z_{N,\nu}=diag(m_{\nu^{\alpha}})$, where $\alpha=1...6$. Here, $M_{\nu}$ is the mass matrix of neutrinos, and $diag(m_{\nu^{\alpha}})$ is the diagonalized mass matrix. We use the matrix $Z_{N,\nu}^{T}$ in the leading of $\zeta$, which is defined as $\zeta=\frac{v_{u}}{{v}_{\eta}}(Y_{\nu}^{T})_{3\times 3}\cdot(Y_{X})_{3\times 3}^{-1}$. Since the neutrino Yukawa coupling $Y_{\nu}$ is very small, all elements of $\zeta$ are very small $\zeta_{IJ}\ll 1$. In the leading-order approximation of $\zeta$, $Z_{N,\nu}^{T}$ can be expressed as22sm10
$Z_{N_{f}}^{T}=\left(\begin{array}[]{cc}\mathcal{S}^{T}&0\\ 0&\mathcal{R}^{T}\end{array}\right)\left(\begin{array}[]{cc}1-\frac{1}{2}\zeta^{\dagger}\zeta&-\zeta^{\dagger}\\ \zeta&1-\frac{1}{2}\zeta\zeta^{\dagger}\end{array}\right).$ (54)
Here, $\zeta$ and $\mathcal{R}$ are matrices used to diagonalize $M_{\nu}^{seesaw}$ and $\sqrt{2}{v}_{\eta}Y_{X}$.
$\zeta^{T}M_{\nu}^{seesaw}\zeta=diag(m_{\nu 1},m_{\nu 2},m_{\nu 3}),$
@@
$M_{\nu}^{seesaw}$ is expressed as
@@
The neutrino mass mixing matrix is the sum of the tree-level result and the one-loop level correction. It is expressed as follows(I,J=1,2,3)
@@
Clearly, the matrix $M^{sum}$ includes a loop correction that also has a seesaw structure. At a loop level, we obtain a corrected effective light neutrino mass matrix of the form22sm10
@@
$\mathcal{H}=(\mathcal{M}_{\nu}^{eff})^{\dagger}\mathcal{M}_{\nu}^{eff}$ (61)
@@
$p=\sqrt{a^{2}-3b},~{}~{}~{}\phi=\frac{1}{3}\arccos\Big{(}\frac{1}{p^{3}}(a^{3}-\frac{9}{2}ab+\frac{27}{2}c)\Big{)},~{}~{}~{}a=\mathrm{Tr}(\mathcal{H})$
@@
$m_{\nu 1}<m_{\nu 2}<m_{\nu 3},~{}~{}~{}m_{\nu 3}^{2}=m_{1}^{2},~{}~{}~{}m_{\nu 1}^{2}=m_{2}^{2},~{}~{}~{}m_{\nu 2}^{2}=m_{3}^{2},$
@@
The neutrino mass spectrum with inverse order (IO) is as follows
$m_{\nu 1}>m_{\nu 2}>m_{\nu 3},~{}~{}~{}m_{\nu 3}^{2}=m_{1}^{2},~{}~{}~{}m_{\nu 1}^{2}=m_{2}^{2},~{}~{}~{}m_{\nu 2}^{2}=m_{3}^{2},$
$\Delta m^{2}_{\odot}=m_{\nu 2}^{2}-m_{\nu 1}^{2}=p(\cos\phi-{\frac{1}{\sqrt{3}}}\sin\phi)<0,$

### v2

@@ I Introduction
Recent experimental data, such as neutrino oscillation experiments released by the Daya Bay Laboratory in 2024DayaBay1 , have further promoted studies on neutrino masses and lepton flavor violating processesDayaBay2 ; DayaBay3 ; lep1 . These data not only provide precise measurements of neutrino mixing angles but also impose new constraints on the upper and lower bounds of neutrino masses. Such constraints are vital for unraveling the nature of neutrinos and their role in the universe. The present $3\sigma$ limits for the neutrino expriment data are6pdg
@@
$6.99\times 10^{-5}\leq\bigtriangleup{m^{2}}_{\odot}\leq 8.07\times 10^{-5}$,
@@
Based on the MSSM, the minimal supersymmetric extension of the SM with local B-L gauge symmetry (N-B-LSSM) is obtained by extending the gauge symmetry group to $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}\times U(1)_{B-L}$7h . The N-B-LSSM introduces three additional Higgs singlets and three generation right-handed neutrinos on the basis of the MSSM. It generates the tiny mass of light neutrinos through the seesaw mechanism. And it provides an effective solution to the $\mu$ problem that the MSSM cannot resolve. In the N-B-LSSM, the Higgs singlet $S$ acquires a non-zero vacuum expectation value (VEV) $\frac{v_{S}}{\sqrt{2}}$, and its superpotential $\lambda SH_{u}H_{d}$ induces an effective $\mu$ term $\mu=\frac{\lambda v_{S}}{\sqrt{2}}$. This mechanism replaces the $\mu H_{u}H_{d}$ term in the MSSM and alleviates the $\mu$ problem. With the introduction of three Higgs singlets, the neutral CP-even Higgs mass matrix is extended to a $5\times 5$ dimension, providing a more natural explanatory framework for the 125 GeV Higgs mass. This work assumes R-parity conservation. In the N-B-LSSM, R-parity is automatically conserved (without the need for manual addition), as the conservation law is defined as: $R_{p}=(-1)^{3(B-L)+2S}$. Other terms in superpotential are allowed. For convenience, they are not considered in this study, because their effects can be combined in the already present terms, and they do not have a substantial new impact on our research.
In the N-B-LSSM, we have discussed certain two-loop contributions to muon anomalous magnetic dipole moment, lepton flavour-violating decays $l_{j}\rightarrow l_{i\gamma},l_{j}\rightarrow 3l_{i}$ and $\mu\rightarrow e+q\overline{q}$, but the neutrino mass problem has not yet been studied7h . In this paper, we study the neutrino mass correction problem under the N-B-LSSM, neutrinos acquire tiny masses at tree level through the seesaw mechanism8seesaw ; 9sm7 ; 102zlxzqqb ; 11Minkowski , facilitated by the introduction of right-handed neutrino fields. However, the one-loop corrections are crucial not only for neutrino mass and mixing10yyl22 ; 11sm but also significantly impact other physical processes. Loop corrections play essential roles in lepton flavor violating processes such as the radiative decays $l_{j}\rightarrow l_{i}\gamma~{}(i,~{}j=1,~{}2,~{}3,~{}i\neq j)$lep1 with zero tree-level contribution. Neutrinos have tiny masses, which implies lepton flavor violation. Therefore, we take into account the important one-loop corrections in this work. Our corrections arise from various sources, including virtual slepton-chargino, sneutrino-neutralino, and Higgs-charged lepton loops. The inclusion of one-loop effects is essential to match the precise experimental data on neutrino oscillations, which provide stringent constraints on the neutrino mass squared differences and mixing angles.
In Ref.12tff , the authors discuss the supersymmetric neutrino mass under the one-loop approximation using the on-shell renormalization scheme without taking into account lepton number conservation and R-parity. There are also some work related to the mass and mixing of neutrinos13yyl ; 101zlxz ; 14yyl33 ; 15tff8 . This paper focuses on the one-loop corrections to the neutrino mass matrix in the N-B-LSSM. We employ the method of mass eigenstates to calculate these corrections and analyze their impact on the neutrino mass spectrum and mixing angles. Our study aims to explore the parameter space of the N-B-LSSM that is consistent with the current neutrino oscillation data. By doing so, we aim to shed light on the underlying mechanisms of neutrino mass generation and to provide insights into the new physics beyond the SM.
@@ II The N-B-LSSM
Under this model, $Y_{u,d,e,\nu,\chi}$ in the superpotential denote the Yukawa coupling coefficients. While $\lambda$, $\lambda_{2}$ and $\kappa$ are dimensionless couplings constant. $\hat{\chi}_{1},~{}\hat{\chi}_{2},~{}\hat{S}$ are three Higgs singlets.
@@
The specific explicit expressions of the neutral components of the two Higgs doublets and three Higgs singlets are as follows
@@ III ANALYTICAL FORMULA
In this section, we compute the one-loop correction to the neutrino mass under the N-B-LSSM. We perform the chiral decomposition of the fermion propagator in loop diagrams, where $P_{L}=\frac{1-\gamma_{5}}{2}$ and $P_{R}=\frac{1+\gamma_{5}}{2}$ present chiral projection operator. The general form of the self-energy of $\nu_{i}-\nu_{j}$ is as follows
$S_{ij}(k)={a_{ij}m_{j}P_{L}}+{b_{ij}m_{i}P_{R}}+{c_{ij}{k\!\!\!/}P_{L}}+{d_{ij}{k\!\!\!/}P_{R}}.$ (31)
Considering the one-loop correction, $a_{ij},b_{ij},c_{ij}$ and $d_{ij}$ are functions of $k^{2}$. In the one-loop correction calculation, $k^{2}$ is the external neutrino momentum square, which is small. So, we perform a Taylor expansion of $a_{ij},b_{ij},c_{ij}$ and $d_{ij}$ according to $k^{2}$. Taking into account the leading-order and next-to-leading order contributions of $a_{ij},b_{ij},c_{ij}$ and $d_{ij}$, we obtain12tff ; 21tff12
@@
$S_{ij}^{REN}(k)u_{j}(k)|_{k^{2}=m_{j}^{2}}=0,$
$\bar{u}_{i}(k)S_{ij}^{REN}(k)|_{k^{2}=m_{i}^{2}}=0.$ (34)
From Eqs. (11)(12)(13), we can get the following solution
@@
$S_{ij}^{REN}(k)=({m_{i}}^{2}b_{ij}^{1}+{m_{i}}^{2}c_{ij}^{1}+{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+a_{ij}^{1}k^{2})m_{j}P_{L}$
$\hskip 42.67912pt+({m_{j}}^{2}a_{ij}^{1}+{m_{j}}^{2}c_{ij}^{1}+{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+b_{ij}^{1}k^{2})m_{i}P_{R}$
$\hskip 42.67912pt+\Big{(}-{m_{j}}^{2}a_{ij}^{1}-{m_{i}}^{2}b_{ij}-({m_{i}}^{2}+{m_{j}}^{2})c_{ij}^{1}-{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+b_{ij}^{1}k^{2}\Big{)}k\!\!\!/P_{L}$
$\hskip 42.67912pt+\Big{(}-{m_{i}}^{2}{m_{j}}^{2}a_{ij}^{1}-{m_{i}}^{2}{m_{j}}^{2}b_{ij}^{1}-{m_{i}}^{2}{m_{j}}^{2}c_{ij}^{1}-({m_{i}}^{2}+{m_{j}}^{2})f_{ij}^{1}+d_{ij}^{1}k^{2}\Big{)}k\!\!\!/P_{R}$
$\hskip 42.67912pt=(k\!\!\!/-m_{j})\hat{S}_{ij}(k)(k\!\!\!/-m_{i}),$ (36)
@@
$\hat{S}_{ij}(k)=a_{ij}^{1}m_{j}P_{R}+b_{ij}^{1}m_{i}P_{L}+c_{ij}^{1}(m_{i}P_{L}+m_{j}P_{R}+k\!\!\!/P_{R})$
$\hskip 42.67912pt+d_{ij}^{1}(m_{i}P_{R}+m_{j}P_{R}+k\!\!\!/P_{L}).$ (37)
@@
$\delta{m_{ij}^{R}}=({m_{j}}^{2}a_{ij}^{1}+{m_{j}}^{2}c_{ij}^{1}+{m_{i}}^{2}{m_{j}}^{2}d_{ij}^{1}+b_{ij}^{1}k^{2})m_{i},$ (38)
where $\delta{Z_{ij}^{L}}(\delta{Z_{ij}^{R}})$ is left-handed(right-handed) wave function renormalization constant. $\delta{m_{ij}^{L}}(\delta{m_{ij}^{R}})$ is left-handed(right-handed) mass renormalization constant. They are used to construct the renormalization propagator.
@@
$=(\delta_{ij}+\delta Z_{ij}^{L})(k\!\!\!/-m_{i}^{tr}-\delta m_{ij}^{L}+\delta Z_{ij}^{L}m_{i}^{tr})P_{L}$
$+(\delta_{ij}+\delta Z_{ij}^{R})(k\!\!\!/-m_{j}^{tr}-\delta m_{ij}^{R}+\delta Z_{ij}^{R}m_{j}^{tr})P_{R},$ (39)
@@
$\delta m^{loop}_{ij}=3m_{i}^{tr}(m_{i}^{tr})^{2}a_{ij}^{1}+\Big{(}m_{i}^{tr}m_{j}^{tr}+(m_{i}^{tr})^{2}+(m_{j}^{tr})^{2}\Big{)}m_{i}^{tr}b_{ij}^{1}$
$\hskip 34.14322pt+\Big{(}(m_{i}^{tr})^{2}m_{j}^{tr}+3m_{i}^{tr}(m_{j}^{tr})^{2}\Big{)}c_{ij}^{1}+\Big{(}3(m_{i}^{tr})^{2}m_{j}^{tr}+m_{i}^{tr}(m_{i}^{tr})^{2}\Big{)}d_{ij}^{1}.$ (40)
@@
$\theta_{w}$ denotes the Weinberg angle, and $\theta^{\prime}_{w}$ denotes $B-L$ mixing angle. $U^{V}$ denotes the rotation matrix to diagonalize the neutrino mass matrix.
@@
$Z^{E}$ denotes the rotation matrix to diagonalize the slepton mass squared matrix. We introduce them in detail in appendix A.
@@
The masses of neutrinos are obtained through the rotation matrix $U^{V}$ using the formula $(U^{V})^{T}M_{\nu}U^{V}=diag(m_{\nu^{\alpha}})$, where $\alpha=1...6$. Here, $M_{\nu}$ is the mass matrix of neutrinos, and $diag(m_{\nu^{\alpha}})$ is the diagonalized mass matrix. We use the matrix $(U^{V})^{T}$ in the leading order of $\zeta=\frac{v_{u}}{{v}_{\eta}}(Y_{\nu}^{T})_{3\times 3}\cdot(Y_{X})_{3\times 3}^{-1}$. Since the neutrino Yukawa coupling $Y_{\nu}$ is very small, all elements of $\zeta$ are much smaller than 1 ($\zeta_{IJ}\ll 1,~{}I,~{}J=1,~{}2,~{}3$). In the leading-order approximation of $\zeta$, $(U^{V})^{T}$ can be expressed as
$(U^{V})^{T}=\left(\begin{array}[]{cc}\mathcal{S}^{T}&0\\ 0&\mathcal{R}^{T}\end{array}\right)\left(\begin{array}[]{cc}1-\frac{1}{2}\zeta^{\dagger}\zeta&-\zeta^{\dagger}\\ \zeta&1-\frac{1}{2}\zeta\zeta^{\dagger}\end{array}\right).$ (54)
Here, $\mathcal{S}$ and $\mathcal{R}$ are matrices used to diagonalize $M_{\nu}^{seesaw}$ and $\sqrt{2}{v}_{\eta}Y_{X}$.
$\mathcal{S}^{T}M_{\nu}^{seesaw}\mathcal{S}=diag(m_{\nu 1},m_{\nu 2},m_{\nu 3}),$
@@
Here to make the discussion simple, we assume that $Y_{X}$ is the diagonal matrix and $\mathcal{R}=1$. $M_{\nu}^{seesaw}$ is expressed as
@@
The neutrino mass mixing matrix is the sum of the tree-level result and the one-loop correction. It is expressed as follows(I, J=1, 2, 3)
@@
Clearly, the matrix $M^{sum}$ includes one-loop correction that also has a seesaw structure. At one loop level, we obtain a corrected effective light neutrino mass matrix of the form22sm10
@@
$\mathcal{H}=(\mathcal{M}_{\nu}^{eff})^{\dagger}\mathcal{M}_{\nu}^{eff}.$ (61)
@@
$p=\sqrt{a^{2}-3b},~{}~{}~{}\phi=\frac{1}{3}\arccos\Big{(}\frac{1}{p^{3}}(a^{3}-\frac{9}{2}ab+\frac{27}{2}c)\Big{)},~{}~{}~{}a=\mathrm{Tr}(\mathcal{H}),$
@@
$m_{\nu 1}<m_{\nu 2}<m_{\nu 3},~{}~{}~{}m_{\nu 1}^{2}=m_{1}^{2},~{}~{}~{}m_{\nu 2}^{2}=m_{2}^{2},~{}~{}~{}m_{\nu 3}^{2}=m_{3}^{2},$
$\Delta m^{2}_{\odot}=m_{\nu 2}^{2}-m_{\nu 1}^{2}=\frac{2}{\sqrt{3}}p\sin\phi>0,$
$\Delta m^{2}_{A}=m_{\nu 3}^{2}-m_{\nu 2}^{2}=p(\cos\phi+{\frac{1}{\sqrt{3}}}\sin\phi)>0.$ (64)
The neutrino mass spectrum with inverse order (IO) is as follows
$m_{\nu 3}<m_{\nu 1}<m_{\nu 2},~{}~{}~{}m_{\nu 3}^{2}=m_{1}^{2},~{}~{}~{}m_{\nu 1}^{2}=m_{2}^{2},~{}~{}~{}m_{\nu 2}^{2}=m_{3}^{2},$